OpenStudy (anonymous):

I need help on two math questions!!!! Directions: Find the points of intersection of the equations. 1) y=x^3 y=x 2) y=x^3-4 y=-(x-2)

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

at the point of intersection you have y1=y2 \[x^3=x \rightarrow x^3-x=0\rightarrow x(x^2-1)=0 \rightarrow x=0,-1,+1\]

OpenStudy (anonymous):

use the same logic for the second one

OpenStudy (anonymous):

Can you kind of show me how to do it?

OpenStudy (dan815):

hey

OpenStudy (anonymous):

Hey

OpenStudy (dan815):

1) y=x^3 y=x lets try to get a picture of whats going on

OpenStudy (dan815):